OpenStudy (anonymous):

Multiply the quantity of 3 plus 5 square root 2 times the quantity of 4 minus square root 2

OpenStudy (anonymous):

\[\left( 3 + 5\sqrt{2} \right) \left( 4 - \sqrt{2} \right)\]

OpenStudy (anonymous):

FOIL it out... F: 3*4 O: (3)(-\(\sqrt{2}) \) I: (\(5\sqrt{2}\) )(4) L: (\( 5\sqrt{2})(-\sqrt{2})\) multiply time out...

OpenStudy (anonymous):

F: 12 O: -3sqrt2 I: 20sqrt2 L: -10

OpenStudy (anonymous):

can you take it from here?

OpenStudy (anonymous):